OpenStudy (anonymous):

An LCO Flat Screen TV at Best Buy increased in price from $900 to $1,200. What was the percent of increase?

OpenStudy (anonymous):

if base price is 900 then it is 33.3%

OpenStudy (anonymous):

300/900= 3/9 =1/3 * 100 = 33.3%